• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 00:05:39 on October 27, 2009

    Entry number 296844

    keyword=Swing Shift Summary

    10/26/09 Swing Shift Summary

    On shift:
    SL: Ellie Long
    TO: Katherine Myers
    3rd: Salvatore Frullani

    · 16:00 Good evening, Hall A! Starting shift at the end of an access, sweep in progress. It looks like the Left HRS is up and we'll be using it tonight.


    · 17:25 We are now in Beam Permit following two sweeps (after the first was completed, they accidentally put us back in Restricted Access and had to do another sweep). MCC sending tuned beam for quick testing.


    · 18:23 Harp scan done. Moved to aluminum dummy for spot++ check.


    · 18:36 The raster came out small. Called MCC who is now checking the settings on the raster magnets.


    · 18:57 MCC is having problems getting the raster magnet up to 10x10 (MCC Units). The voltages are not reaching what they were in the past. Expert called. Beam set to 1uA.


    · 19:43 MCC has done what they can with raster. Switching to Carbon Hole target to do spot++ on 10uA.


    · 20:04 Raster is better, but still fuzzy. At this point, we're going to run with it.


    · 21:51 TO noticed that the Target Alarm Handler was not on and, upon trying to open it, found the file to be missing. Expert was called, it is now on again.


    · 00:02 Shift ends with left Q2 tripping.